OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

45x+28.50y=4770 x+y=150

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

well those are the equations, then you need to solve for x or y in one of the equations and plug it into the other equation

OpenStudy (anonymous):

thats what ineed help with

OpenStudy (anonymous):

x=150-y so for x in the other equation, plug in 150-y and then solve for y